Assume that a sample is used to estimate a population proportion p. Find the margin of error E that corresponds to the given statistics and confidence level. Round the margin of error to four decimal places.

95% confidence; n = 356, x = 49
Answer #1

Ans) 0.0358

Explanation:

Sample \: proportion\: \: \hat{p} = \frac{x}{n} = \frac{49}{356} = 0.138

Critical value for 95% confidence =

z_{0.05/2} = 1.96

/* we can find it using excel function: =NORM.S.INV(0.05/2) */

Margin\: of\: error\: E = z_{\alpha/2}\sqrt{\frac{\hat{p}(1-\hat{p})}{n}}

=1.96\sqrt{\frac{0.138(1-0.138)}{356}}

=0.0358 ans,

/* PLEASE UPVOTE */